5(2x-1)-4x=6x+6 Solve the equation
I Get 11 when i try but that doesn't work when you plug it back into the equation.

Answer by venugopalramana(3286)

5(2x-1)-4x=6x+6 Solve the equation
5*2x-5*1-4x=6x+6
10x-5-4x-6x=6
x(10-4-6)=6+5=11
0*x=11..or 0=11 which is wrong..so the problem is not proper that is why you did not get the confirmation on verification ..rewite the problem as
5(2x-1)-3x=6x+6..and you will get the correct answer..